A 0.1m long elastic band extends 5mm when a load of 80N is hung from its end. Calculate the strain on the band
1
Expert's answer
2019-04-05T10:28:26-0400

We can find the strain on the band from the formula:


"\epsilon = \dfrac{\Delta l}{l_0} = \dfrac{l - l_0}{l_0},"

here, "\Delta l" is the extension of the band (or change in the length of the band), "l_0" is the initial length of a band being stretched and "l" is the length of the band after it has been stretched.

Then, we get:


"\epsilon = \dfrac{(0.1 m + 0.005 m) - 0.1 m}{0.1 m} = 0.05."

Answer:

"\epsilon = 0.05."